OpenStudy (anonymous):

Fan and medal? Question below!! Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - y = -11 2x + 3y = 5 -4x - 9y = -19 -10y = -30 4x + 3y = 5 2y = -6 7x - 3y = -11 9x = -6 12x - 3y = -33 14x = -28

OpenStudy (adajiamcneal):

I would suggest first solving the system of equations 4x - y = -11 2x + 3y = 5 for x and y...

OpenStudy (adajiamcneal):

here is what I did (I used substitution method) 4x - y = -11 --> y = 4x + 11 2x + 3y = 5 = 2x + 3(4x + 11) = 5 distribute 2x + 12x + 33 = 5 add like terms and subtract both sides by 33 14x = -28

OpenStudy (anonymous):

Thanks beautiful :)

OpenStudy (adajiamcneal):

anytime :)